OpenStudy (anonymous):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits.

OpenStudy (shamil98):

\[\huge Area~of~a~\square = s^2\] so\[\huge 6.25^2 = area\]

OpenStudy (shamil98):

\[\huge area = 6.25 \times 6.25\] if you're wondering what 6.25^2 is..

OpenStudy (anonymous):

Soo it would be 39.0625?

OpenStudy (shamil98):

Yes!